# When Single-Agent with Skills Replace Multi-Agent Systems and When They Fail

### v1

@@ Abstract
(a) From multi-agent coordination to a single agent with skills.
@@
Figure 1: Skill-based agents: efficiency gains and scaling limits. (a) Compiling multi-agent systems into single-agent skill libraries reduces communication overhead, cutting latency and token usage. (b) Skill selection accuracy degrades non-linearly as libraries grow, exhibiting a phase transition at a capacity threshold. As skill libraries grow, the increased size and semantic confusability among skills drive this degradation; hierarchical routing restores reliable selection by organizing skills into structured categories. Visualization modified from Nana Banana.
@@ 1 Introduction
To answer this question, we study the scaling behavior of skill selection in LLMs. Drawing on principles from cognitive science, we hypothesize that skill selection exhibits capacity-limited scaling analogous to human decision-making. Our experiments confirm this hypothesis. Across controlled skill libraries ranging from 5 to 200 skills, we find that selection accuracy degrades non-linearly, following a phase transition pattern: accuracy remains high when library size is below a critical threshold, then drops sharply beyond this capacity. This degradation is also driven by semantic confusability among skills, rather than by library size alone. This finding connects LLM behavior to similarity-based interference in human memory retrieval (Shepard, 1987; Anderson, 1974). We further show that hierarchical routing: decomposing skill selection into coarse-to-fine decisions can effectively mitigate this degradation when flat selection fails when skill set scales up. This mirrors findings from cognitive science on chunking (Chase and Simon, 1973) and menu design (Miller, 1981).
@@ 2.3.4 Optimization Objective
The final Single-Agent System is thus equipped with the compiled skill set:
@@ 4.1 Cognitive Foundations
Our scaling hypothesis draws on established principles from cognitive psychology and decision science. Three foundational theories are particularly relevant:
@@ 4.3 The Scaling Law
where $\alpha\leq 1$ is asymptotic accuracy at small $|\mathbf{S}|$, $\kappa$ is the capacity threshold, $\gamma>1$ controls the sharpness of the phase transition, and $\epsilon>0$ represents sensitivity to semantic interference.
@@ Results.
Figure 2 presents selection accuracy as a function of skill library size across both evaluated models and demonstrates the quality of the scaling law fit. Both models exhibit a sharp decline in selection accuracy as library size increases. At small scales ($|\mathbf{S}|\leq 20$), accuracy remains above 95%, but degrades rapidly beyond $|\mathbf{S}|=50$, falling to approximately 20% at $|\mathbf{S}|=200$. The results support our hypothesis of non-linear phase transition. The accuracy curves exhibit a characteristic sigmoid decay, with the steepest decline occurring near $|\mathbf{S}|\approx 50$100100, aligning with the estimated $\kappa$ values. The decay rate $\gamma>1$ (1.71 and 1.65) indicates faster-than-linear degradation beyond the capacity threshold, consistent with predictions from cognitive load theory and Hick’s Law from cognitive psychology, which predicts logarithmic increases in decision time (and by extension, error rates) with the number of choices. However, we expect to see that stronger models with larger $\kappa$ but the fitted parameters contradict with this. We note the fitting itself contains error and the more models with different capcities need to be tested. There a slight increase in accuracy as $|\mathbf{S}|$ grows from 5 to approximately 20. This likely reflects improved task-skill coverage: at very small library sizes, some tasks lack well-matched skills, while moderate expansion improves coverage before selection overload begins. This non-monotonic behavior at small $|\mathbf{S}|$ does not affect our main findings, which concern the capacity-limited regime at $|\mathbf{S}|\gg 20$.
@@ 5.6 Practical Guidelines and Limitation Discussion
Design Guidelines Based on our findings, we recommend the following guidelines for skill library design in LLM-based agents: 1. Monitor library size: Track $|\mathbf{S}|$ relative to the model’s capacity ($\kappa\approx 50$100100 for the tested GPT models). Performance degrades sharply beyond this threshold. 2. Minimize confusability: Before adding skills, audit semantic overlap. Merge or differentiate skills with similarity above threshold rather than accumulating near-duplicates. 3. Adopt hierarchy at scale: For large $|\mathbf{S}|\gg\kappa$, implement hierarchical routing with confusability-aware grouping. Each stage should involve $\kappa$ options. 4. Invest in descriptors: Since selection relies on descriptors (H2), invest effort in crafting or optimizing distinctive, specific descriptions. 5. Match model to task: Stronger models show higher $\kappa$ and better confusability resistance. For applications with inherently large or confusable skill sets, model capability investment yields direct accuracy benefits. 6. Consider alternative architectures: For large skill sets, the scaling limitations suggest that multi-agent architectures with specialized routers may outperform monolithic single-agent approaches.

### v2

@@ Abstract
(a) From multi-agent systems to a single agent with skills.
@@
Figure 1: Skill-based agents: efficiency gains and scaling limits. (a) Compiling multi-agent systems into single-agent skill libraries reduces communication overhead, cutting latency and token usage. (b) Skill selection accuracy degrades non-linearly as libraries grow, exhibiting a phase transition at a capacity threshold. As skill libraries grow, the increased size and semantic confusability among skills drive this degradation; hierarchical routing restores reliable selection by organizing skills into structured categories.
@@ 1 Introduction
To answer this question, we study the scaling behavior of skill selection in LLMs. Drawing on principles from cognitive science, we hypothesize that skill selection exhibits capacity-limited scaling analogous to human decision-making. Our experiments confirm this hypothesis. Across controlled skill libraries ranging from 5 to 200 skills, we find that selection accuracy degrades non-linearly, following a phase transition pattern: accuracy remains high when library size is below a critical threshold, then drops sharply beyond this capacity. This degradation is also driven by semantic confusability among skills, rather than by library size alone. This finding connects LLM behavior to similarity-based interference in human memory retrieval (Shepard, 1987; Anderson, 1974). We further show that hierarchical routing: decomposing skill selection into coarse-to-fine decisions can effectively mitigate this degradation when flat selection fails when skill library scales up. This mirrors findings from cognitive science on chunking (Chase and Simon, 1973) and menu design (Miller, 1981).
@@ 2.3.4 Optimization Objective
The final Single-Agent System is thus equipped with the compiled skill library:
@@ 4.1 Cognitive Foundations
Our scaling hypothesis draws on established principles from cognitive psychology and decision science. Four foundational theories are particularly relevant:
@@ 4.3 The Scaling Law
where $\alpha\leq 1$ is asymptotic accuracy at small $|\mathbf{S}|$, $\kappa$ is the capacity threshold when accuracy drops to half, $\gamma>1$ controls the sharpness of the phase transition, and $\epsilon>0$ represents sensitivity to semantic interference.
@@ Results.
Figure 2 presents selection accuracy as a function of skill library size across both evaluated models. We evaluate at $|\mathbf{S}|\in\{5,10,20,25,30,40,50,75,100,150,200\}$ and fit the proposed scaling law (Eq. 10). Both models exhibit substantial accuracy degradation as library size increases. At small scales ($|\mathbf{S}|\leq 20$), accuracy remains above 90%, but degrades steadily beyond $|\mathbf{S}|=30$, falling to approximately 20% at $|\mathbf{S}|=200$. We observe a slight increase in accuracy as $|\mathbf{S}|$ grows from 5 to approximately 20. This likely reflects improved task-skill coverage: at very small library sizes, some tasks may lack well-matched skills, while moderate expansion improves coverage before selection overload begins. This non-monotonic behavior at small $|\mathbf{S}|$ does not affect our main findings, which concern the capacity-limited regime at larger library sizes. The decay exponent $\gamma>1$ (1.72 and 1.56 respectively) indicates super-linear degradation, consistent with phase transition behavior rather than gradual decay.
The fitted capacity parameter $\kappa$ is 91.8 for GPT-4o-mini and 83.5 for GPT-4o, representing the library size at which accuracy drops to half its maximum ($\alpha/2\approx 0.48$$0.49$). Interestingly, GPT-4o exhibits a slightly lower $\kappa$ than GPT-4o-mini, contrary to the intuition that stronger models should have higher capacity. We offer two possible explanations: (1) this difference may reflect fitting variance given limited data points, or (2) model capability and selection capacity may be partially independent; a more capable model is not necessarily better at discriminating among many similar options. Further experiments across diverse model families are needed to clarify this relationship. We note that $\kappa$ should not be interpreted as the “onset” of degradation. The inflection point, where accuracy begins to decline most rapidly, occurs earlier. For practical applications, we recommend keeping flat skill libraries below this inflection point, or adopting hierarchical routing for larger libraries.
@@ 5.6 Practical Guidelines and Limitation Discussion
Design Guidelines Based on our findings, we recommend the following guidelines for skill library design in LLM-based agents: 1. Monitor library size: Track $|\mathbf{S}|$ relative to the model’s capacity ($\kappa\approx 50$100100 for the tested GPT models). Performance degrades sharply beyond this threshold. 2. Minimize confusability: Before adding skills, audit semantic overlap. Merge or differentiate skills with similarity above threshold rather than accumulating near-duplicates. 3. Adopt hierarchy at scale: For large $|\mathbf{S}|\gg\kappa$, implement hierarchical routing with confusability-aware grouping. Each stage should involve $\kappa$ options. 4. Invest in descriptors: Since selection relies on descriptors (H2), invest effort in crafting or optimizing distinctive, specific descriptions. 5. Match model to task: Stronger models show higher $\kappa$ and better confusability resistance. For applications with inherently large or confusable skill libraries, model capability investment yields direct accuracy benefits. 6. Consider alternative architectures: For large skill libraries, the scaling limitations suggest that multi-agent architectures with specialized routers may outperform monolithic single-agent approaches.